OpenStudy (anonymous):

Simplify 2-x/3 ------ 5+x/8

OpenStudy (anonymous):

Openstudy is a place to study and learn. Try to apply what you just learned rather than having the website do your homework.

OpenStudy (anonymous):

multiply top and bottom by 24 to clear the fractions. don't forget distributive law

OpenStudy (anonymous):

got it or no?

OpenStudy (anonymous):

\[\frac{2-\frac{x}{3}}{5+\frac{x}{8}}\times \frac{24}{24}\]

OpenStudy (anonymous):

2/5?

OpenStudy (anonymous):

oh no dear your answer will have an x in it. lets go slow

OpenStudy (anonymous):

only thing we are trying to do is to clear the fractions

OpenStudy (anonymous):

so least common multiple of 3 and 8 is just 3*8=24

OpenStudy (anonymous):

multiply top by 24 you get \[(2-\frac{x}{3})\times 24=2\times 24-\frac{x}{3}\times 24=48-8x\]

OpenStudy (anonymous):

so numerator is \[48-8x\]

OpenStudy (anonymous):

denominator is similar \[(5+\frac{x}{8})\times 24=5\times 24+\frac{x}{8}\times 24=120+3x\]

OpenStudy (anonymous):

120+3x would be the denominator?

OpenStudy (anonymous):

yes you got it!

OpenStudy (anonymous):

Wow thank you so much!

OpenStudy (anonymous):

so "final answer is ..."

OpenStudy (anonymous):

\[\color{red}{\frac{48-8x}{120+3x}}\] !